OpenStudy (anonymous):

given the midpoint of a segment is (3,7) and one endpoint has the coordinates (10,2) what is the coordinates of the other endpoint?

OpenStudy (anonymous):

The x-coordinate of the other endpoint is the solution to this equation: \[\frac{ 10+x }{ 2 }=3\] because the formula for the midpoint is the average of the two endpoint x-values. So if you add the two endpoint x-values, then divide by 2, you'll get the midpoint x-value. The same goes for the y-value, except it will be this equation: \[\frac{ 2+y }{ 2 }=7\] Do you know how to solve equations like that?